OpenStudy (anonymous):

HELP ? Simplify the sum. State any restrictions on the variables. x-2/x+3 + 10x/x^2-9

OpenStudy (dtan5457):

\[\frac{ x-2 }{ x+3 }\div \frac{ 10x }{ x^2-9 }\]

OpenStudy (dtan5457):

What looks factorable?

OpenStudy (anonymous):

ummmmm.... i dont really understand what you mean @dtan5457

OpenStudy (dtan5457):

To simplify rational expressions, you want to cross out as many terms as possible. Look at x^2-9. What does that simplify to?

OpenStudy (anonymous):

i dont know D: 1?

OpenStudy (anonymous):

sorry math is my worst subject

OpenStudy (dtan5457):

\[a^2-b^2=(a+b)(a-b)\]

OpenStudy (dtan5457):

x^2-9= (x+3)(x-3)

OpenStudy (dtan5457):

Got it?

OpenStudy (anonymous):

yee

OpenStudy (dtan5457):

Ok, now if you remember from 5th grade lol, what do you do when you divide fractions?

OpenStudy (anonymous):

multiply?

jagr2713 (jagr2713):

@chlobohoe please avoid giving direct answers :D

OpenStudy (dtan5457):

You switch the 2nd fraction (reciprocal), then multiply

OpenStudy (anonymous):

ye

OpenStudy (anonymous):

wait what fractions? ahahaha

OpenStudy (dtan5457):

\[\frac{ x-2 }{ x+3 } \times \frac{ (x+3)(x-3) }{ 10x }\]

OpenStudy (dtan5457):

Once you switch it, it should look like that. Now what terms can we cross out?

OpenStudy (anonymous):

ohhh um ..x+3 and 10x?

OpenStudy (anonymous):

dont you suppose to multiply going straight across right? or no?

OpenStudy (dtan5457):

You are, but if there is a same term in the numerator and denominator, you can cross them out.

OpenStudy (dtan5457):

In this case's x+3

OpenStudy (dtan5457):

\[\frac{ (x-2)(x-3) }{ 10 }\]

OpenStudy (dtan5457):

Opps forgot the x on the 10

OpenStudy (anonymous):

yea i understand

OpenStudy (anonymous):

so then do that fraction stay like that?

OpenStudy (dtan5457):

Kk wait a second. I think I'm gonna jump off a building

OpenStudy (anonymous):

hahah what's wrong

OpenStudy (dtan5457):

Your question says to add..I divided.

OpenStudy (anonymous):

ohhh .. i wouldnt have noticed loool ..

OpenStudy (dtan5457):

If you will excuse me for 2 minutes while I redo this problem that will be great

OpenStudy (anonymous):

okiee that's fine

OpenStudy (dtan5457):

Ok got it. Let's do this again...sorry

OpenStudy (dtan5457):

\[\frac{ x-2 }{ x+3 }+\frac{ 10x }{ x^2-9 }\] Factor out x^2-9

OpenStudy (dtan5457):

We said that was x+3(x-3)

OpenStudy (anonymous):

yee

OpenStudy (dtan5457):

\[\frac{ x-2 }{ x+3 }+\frac{ 10x }{ (x-3)(x+3) }\]

OpenStudy (dtan5457):

When we add fractions, they must have the same denominator

OpenStudy (dtan5457):

Therefore, we multiply x-2/x+3 by x-3

OpenStudy (dtan5457):

\[\frac{ (x-3)(x-2) }{ (x+3)(x-3) }+\frac{ 10x }{ (x-3)(x+3) }=\frac{ x^2-6x+6+10x }{ }\]

OpenStudy (dtan5457):

The denominator would still be (x+3)(x-3)

OpenStudy (dtan5457):

\[\frac{ x^2+4x+6 }{ (x+3)(x-3) }\]

OpenStudy (dtan5457):

You can factor the top part as \[\frac{ (x+3)(x+2) }{ (x+3)(x-3) }\]

OpenStudy (dtan5457):

Cross out the x+3 and you get\[\frac{ x+2 }{ x-3 }\]

OpenStudy (anonymous):

ohhhhhhhhh so you cancel out the same thing that's on top i see ..

OpenStudy (anonymous):

THANK YOUUUUUUUUUUUUUU for helping

jagr2713 (jagr2713):

Great explaining @dtan5457

OpenStudy (dtan5457):

Yeah, it's just too bad I did the whole problem wrong by dividing instead of adding at first

OpenStudy (dtan5457):

Do you know what the restrictions are? @rawmoli

OpenStudy (anonymous):

no?

OpenStudy (dtan5457):

Whenever you have rational fractions, usually the only restriction is that the denominator CANNOT be 0.

OpenStudy (dtan5457):

\[\frac{ x-2 }{ x+3 }\]

OpenStudy (dtan5457):

What x value would make this invalid?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

sorry i dont understand aahah ..

OpenStudy (dtan5457):

Denominator can't be 0, not numerator

OpenStudy (anonymous):

ohhhhh

OpenStudy (dtan5457):

So x can't be..?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i think

OpenStudy (dtan5457):

lol take your time

OpenStudy (dtan5457):

x+3=0 x=?

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

x=-3, but that's just one of the restrictions

OpenStudy (dtan5457):

\[\frac{ x-2 }{ x+3 }+\frac{ 10x }{ x^2-9 }\]

OpenStudy (dtan5457):

actually i'm not 100% sure but i'm pretty sure if x=3, that would make the 2nd fraction have a 0 and be restricted

OpenStudy (dtan5457):

@Nnesha

OpenStudy (anonymous):

OH SO I WAS RIGHT WITH -3? yesssssssssssssssssssss

OpenStudy (dtan5457):

Yeah, but there might be one more, but i'm not 100% sure so..

OpenStudy (anonymous):

ohh okiee

OpenStudy (anonymous):

thank you for everything